Can someone please help me finish the tables and calculations.
Thanks in advance
Kinetics Results: Rate Law Rate Constant Activation Energy Frequency Factor
Data Sheet Part A: Dependence of Reaction Rate on Concentration Table 1 H Temperature Time Relative mixture ! (seconds) | rate 1000 Reaction BrOsl (eC) . 1235.6 Avg. time 2. 67.S 3. 08,24 Avg. time 2. 74.2 3. 70.74.O Avg. time 24.o' 4 弘·11 0.2 OD3 0.00 8 100 o.oy | 23.5 3. 29. Avg. time 교 Table 2: Calculations for the [I']. [Bro, & [H'] BrOs Hl [I-I
Table 3: Caleulate the effect of concentration on reaction rate for each reagent (m, n, p); Rate kI (Bros1 H1 IH'1 BrOs1 5.13 7.13 .Complete the rate law equation below Rate k [I] [Bro,] [H]° Table 4: Calculate the rate constant k for each reaction mixture Reaction Rate constant k mixture Average
Part B: Determination of the Activation Energy (E.) and Frequency factor (A) of the Reaction Table 5 Temp of reaction mixture Time Temp Relative rate 1000/t Avg. time (s) oC From part A 1. 535 53 2. Avg. time (s) B445 13,5 с eC 12. 3. Avg. time (s) 1os 3. Avg. time (s) 85 5 37.33t C Avi Use the following equation to calculate the activation energy (Ea) k = Ae-Ea/RT Taking logarithms we get: In k =-RT + lnA
Kinetics Table 6: k at each Temperature: from rate law determined in part A, Complete the following table for the four temperatures. Show calculations with correct units. In k T(K) 1/T (K) Reaction Temperature(0) 5.3'c 25C 29 C Using Excel, 1. Plot a scatter graph of "In k versus 1/T".Be sure to label the axis and title your graph 2. Fit a trend line to your plotted data, and obtain both the equation of the trend line and the R2 value. Submit a printout of this graph along with the complete data table with your datasheets & calculations. 3. 4. Calculate and report the activation energy,E and the frequency factor, A, for this reaction (include units!). Show complete calculations for determining E, and A. 10
Kinetics R2 (Correlation Coefficient) Equation of the line: Calculations of Ea: (J/mol) E (kJ/mol Calculations of A (frequency Factor) A (frequency factor) Part C. Effect of a Catalyst upon Reaction Rate Table 7: Reaction mixture 1 Time (s) T- *From part A 53.74 Plus catalyst By what factor did the rate of reaction increase? katalysedknaalyse,-
Kinetics Results: Rate Law Rate Constant Activation Energy Frequency Factor
Answer #1

p= 2.367

Table 4

Rate constants:

1) k= 4.55 × 10^9

2) k= 4.759 × 10^9

3) k= 4.472 × 10^9

4) k= 4.53 × 10^9

k average = 4.57775 ×10^9

Table 5:

Temp °C average time (seconds) temp K relative rate
25 141.2 298 7.08
5.3 344 278.9 2.9
18.5 170 291.5 5.88
22 85 295 11.764
37.33 41 310.33 24.324

  

Table 6:

Reaction temp °C T K 1/ T k ln k
5.3 278.3 .00359
18.5 291.5 .00343
25 298 .00335
29 302 .00331
37 310 .00322

The values of k are known only for the above 4 temperatures in in table 1.

For further calculations of k at different temperatures other than those mentioned in table 1 one need more data for k calculation at 5.3 18.5 29 and 37 °C. I hope you get my point and for furyher queries feel free to contact and then I can help you further with the calculations.

Thank you